# Cascade-driven opinion dynamics on social networks

### v1

@@ V-B1 Dataset description and experimental setting
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, testing three scenarios: low diffusion ($\theta=0.01$), moderate diffusion ($\theta=0.8$), and maximum diffusion ($\theta=1$), where nodes always reshare posts. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 simulation runs.
V-B2 Opinion evolution under FJC vs FJ
@@
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. In Figure 4, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$). Due to space constraints, we report results for $\theta=0.01$ and $\theta=0.8$ (shown in subpanels (i) and (ii)), and for the polarizing vectors $\bm{u}_{B_{1}(1)}$, $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{V_{>1}}$, and $\bm{u}{V_{>1}}$. The full results are available in the SI Appendix.
@@
We start with panels C of Figure 4, where all nodes have equal susceptibility. Compared to the FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—visible from the shorter arrows relative to the distance to the FJ outcomes (black dots)—especially for $\theta=0.01$ (Ci) and $\theta=0.8$ (Cii). This leads to a wider spread of final opinions among central nodes. In contrast, less central nodes are more influenced and their opinions shift closer to those of central ones. Overall, cascades amplify the influence of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible. The reposting probability $\theta$ modulates this effect—low values ($\theta=0.01$) limit opinion spread, while higher values ($\theta=0.8$) increase it. When susceptibility is inversely proportional to centrality (panels B), central nodes are already stubborn by design in both models, leading to similar outcomes. Nonetheless, cascades still shift the opinions of less central nodes more in the FJC model, as they become increasingly influenced by central ones. Finally, with susceptibility proportional to centrality (panels A), network topology and susceptibility act in opposing directions. Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position. As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model. This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
@@
Figure 4: Polarizing opinions vs. node centrality. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$; (i) $\theta=0.01$; (ii) $\theta=0.8$. In each panel, left plots correspond to $P_{4}$-polarization vectors, right plots to $P_{2}$/$P_{3}$ vectors. Dotted and dashed lines show the average initial and final FJC opinions.
V-B3 Polarization
@@
Figure 5 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions. We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC). Due to space constraints, we show only the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others. Full results are in the SI. We first consider FJC with $\theta=0.8$ and $\theta=1$ (blue and purple bars). When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions. A similar pattern, though less pronounced, appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC. Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ. Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape. An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization. For $\theta=0.01$ (green FJC bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude. In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions. For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged. The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
@@
V-C FJC with real cascades
V-C1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [47], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-B, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772, 693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, due to the large size of the network, even optimized methods could not compute these vectors over the full LSCC. We therefore define two subgraphs:
  • Nepal cut degree: includes only nodes with degree in the range $[150,400]$ (where the node degree spans from 2 to 53718 in the original LSCC), balancing network activity and computational feasibility. We extract the LSCC and retain only the cascades involving its nodes.
  • Nepal top cascades: constructed from nodes involved in the 500 longest cascades using a heuristic for the directed Steiner Tree problem [48] to ensure that all the nodes involved in these cascades are in the subgraph.
Table I summarizes the two subgraphs. The Nepal cut degree graph has more nodes but fewer edges, while Nepal top cascades is denser and includes more cascades (hence making it more likely that assumption (A3) is violated).
TABLE I: Summary of the two real-world Nepal datasets subgraphs used for validation
#nodes #edges #cascades max cascade length
Nepal cut degree 30662 737541 4111 4
Nepal top cascades 26321 4222064 6083 21
Each dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades. We assume that this sequence of reshares drives the opinion dynamics in the real world: each reshare is treated as an expression of opinion. At time $t_{k}$, when node $i_{k}$ shares a post, its followers update their opinions according to Equation (17). This process relies solely on assumption (A1); assumptions (A2) and (A3) are not used. We refer to this approach as real opinion dynamics. It is a data-driven heuristic and cannot be analyzed analytically.
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [35, 49], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE II: Statistics of followers, followed and $\theta$ in both Nepal subgraphs
Nepal cut degree
Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
Followers 1.00 8.00 19.00 24.05 34.00 182.00 -
Followees 1.00 13.00 21.00 24.05 32.00 135.00 -
$\theta$ 0.000 0.000 0.000 0.001 0.000 1.000 8807
Nepal top cascades
Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
Followers 1.0 23.0 59.0 160.4 148.0 11022.0 -
Followees 1.0 53.0 109.0 160.4 204.0 6851.0 -
$\theta$ 0.0000 0.0000 0.0000 0.0156 0.0000 1.0000 681
V-C2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we do not visualize the polarizing vectors as in Figure 4, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 6, FJC closely matches the real dynamics in the Nepal cut degree dataset (the two curves are nearly indistinguishable) and shows reasonable agreement in the Nepal top cascades dataset. This is expected, as longer cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly robust even in this more challenging scenario.
Figure 6: ECDF of the opinion shift of nodes, i.e. final opinion minus initial opinion, in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-C3 Polarization
Figure 7 shows polarization outcomes across different settings. Both cascade-based dynamics—FJC and Real—exhibit very low or even negative polarization, in sharp contrast to the FJ model. This suggests that, in this case, cascades tend to reduce polarization, seemingly contradicting the findings from synthetic networks in Section V-B. However, this can be explained by the fact that, as shown in Table II, most nodes have $\theta$ values close to zero, indicating minimal reposting activity. These results are therefore consistent with our earlier observation: when $\theta$ is small, opinion dynamics are largely inactive, preventing polarization from developing.
@@ VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution. To validate the FJC model, we utilized a real cascade dataset. Our results demonstrate that the FJC model is robust and can approximate final opinions accurately, even when many of the underlying assumptions of the model are violated.

### v2

@@ V-C1 Dataset description and experimental setting
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts, which are kept fixed across all simulations. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios. In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table II). To explore other potential values, we referenced typical ranges established in the literature. Hodas et al. in [53] estimated that the retweet probability for each tweet seen is around 0.15% to 0.4%. Based on this, we modelled a moderate diffusion scenario with $\theta=0.1$, and for maximum diffusion, we used $\theta=0.5$. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 independent simulation runs.
V-C2 Opinion evolution under FJC vs FJ
@@
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. Due to space constraints, we will present just $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, the best local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}-P_{3}$ polarization, respectively. Please refer to the Supplemental Information (SI) for the full results. In Figures 6-8, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
@@
We begin our discussion with Figure 8, which corresponds to the scenario with the highest propagation ($\theta=0.5$), as this setting highlights the effects of cascades on opinion dynamics at a larger scale. Afterwards, we will examine how these trends change when $\theta$ is smaller. Let us first consider panel C, where all nodes share the same susceptibility. Compared to the classical FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—evident from the shorter arrows relative to the distance to the FJ outcomes (black dots). This leads to a wider spread of final opinions among central nodes. In contrast, peripheral nodes are more easily influenced, and their opinions shift closer to those of the central ones. Overall, cascades amplify the role of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible followers. When susceptibility is inversely proportional to centrality (panel B), central nodes are already stubborn by design in both models. Final opinions are mostly similar across all networks, although we can observe slightly smaller opinion shifts in FJC. This reduction is due to the parameter $\theta$, which limits opinion propagation. This trend is further confirmed by the figures in the SI, where we also tested even larger values for this parameter: as $\theta$ increases, peripheral nodes shift more strongly in the FJC model, becoming increasingly influenced by central nodes. Finally, with susceptibility proportional to centrality (panel A), network topology and susceptibility act in opposing directions. Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position. As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model. This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 6 and 7. In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared. Notably, in the Karate network, results remain almost unchanged across different $\theta$ values, as its small size allows most nodes to be reached by the cascade even when propagation is limited. Interestingly, central nodes in the FJC model maintain stable behavior even with tighter propagation. This resilience is an intrinsic property of the model, stemming from the structured, sequential nature of its cascades. In fact, highly connected nodes are typically exposed to content early, when active sources are few, making them less susceptible to external influence.
$\bm{\theta=0.01}$
@@
Figure 6: Polarizing opinions vs. node centrality with $\theta=0.01$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
$\bm{\theta=0.1}$
Figure 7: Polarizing opinions vs. node centrality with $\theta=0.1$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
$\bm{\theta=0.5}$
Figure 8: Polarizing opinions vs. node centrality with $\theta=0.5$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
V-C3 Polarization
@@
Figure 9 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions. We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC). We show the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others. We first consider FJC with $\theta=0.5$ (purple bars). When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions. A similar, though less pronounced, pattern appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC. An exception arises for $\bm{u}_{{V_{>1}}^{heu}}$ in the BA network, where polarization decreases because the parameter $\theta$ limits opinion propagation, as discussed in the previous section. Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ. Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape. An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization. For $\theta=0.01$ and $\theta=0.1$ (green and blue bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude. In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions. For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged. The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
@@ VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We validated the FJC model, utilising a real cascade dataset. Our results demonstrate that the FJC model can reliably approximate final opinions accurately, even when many of the underlying assumptions of the model are relaxed or only partially met. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution.